# AlphaApollo: Orchestrating Foundation Models and Professional Tools into a Self-Evolving System for Deep Agentic Reasoning

### v1

@@ 3 Experiments
In this section, we evaluate the AlphaApollo on complex reasoning tasks. We first describe the experimental setup, followed by the main experimental results and detailed analysis.
Table 1: The detailed settings of models in our experiments. Here, output length refers to the maximum response length specified during evaluation; model type characterizes the reasoning paradigm—instruction models directly produce answers that may include explicit multi-step reasoning, while reasoning models first generate extended internal reasoning traces before presenting a structured multi-step solution; and architecture indicates whether the model adopts a Mixture-of-Experts (MoE) or a dense structure.
xxModel Familyxx xxxxModel Namexx Maximum context length Output length Model type Architecture
Qwen2.5 (Yang et al., 2024a) Qwen2.5-7B-Instruct 128K 8k Instruction Dense
Qwen2.5-14B-Instruct 128K 8k Instruction Dense
Qwen2.5-32B-Instruct 128K 8k Instruction Dense
Qwen2.5-72B-Instruct 128K 8k Instruction Dense
Qwen3 (Yang et al., 2025) Qwen3-235B-A22B 80k 32k Reasoning MoE
Llama (Grattafiori et al., 2024) Llama3.3-70B-Instruct 128k 8k Instruction Dense
Experiment settings.
We evaluate AlphaApollo across different model families, including the Qwen2.5 series, Qwen3 series, and Llama series. The model configurations are shown in Table 1. For benchmarks, we use AIME 2024 and AIME 2025, both derived from the American Invitational Mathematics Examination, a competition widely considered one of the most challenging mathematics competitions at the high-school level. These benchmarks evaluate sophisticated problem-solving and reasoning abilities that exceed conventional competition datasets. The two iterations, AIME 2024 and 2025, have distinct problem sets, facilitating annual assessment of robustness and generalization. For the experimental setup, we evaluate all models with sampling parameters of temperature = $0.6$, top-k = $20$, and top-p = $0.95$ for evaluation. Each question is sampled $32$ times to mitigate the impact of randomness. We report both Average@$32$ and Pass@$32$ results: the former reflects the model’s average performance, while the latter indicates the upper bound of the model’s capability.
@@
We summarize the following observations from the experimental results in Table 2. The results demonstrate that AlphaApollo consistently improves performance across different model families and scales. For medium-sized models such as Qwen2.5-14B/32B/72B-Instruct, AlphaApollo achieves substantial gains in both Average@$32$ and Pass@$32$, with improvements of up to $9.16$ % and $23.34$ %, respectively. Notably, for the large model Qwen3-235B-A22B, although the base model already performs strongly, AlphaApollo still provides consistent improvements, showing that our method scales effectively. Moreover, on Llama3.3-70B-Instruct, AlphaApollo delivers remarkable boosts—especially on AIME 2025, where Average@$32$ increases by $16.67\%$ and Pass@$32$ doubles from $23.33\%$ to $46.67\%$. Overall, these results indicate that AlphaApollo can significantly strengthen both the average performance and the capability ceiling of state-of-the-art FMs.
Table 2: Main experiment results (in %). The boldface numbers denote the best results. We also report the absolute improvement of AlphaApollo over the base model for each dataset.
xxBase Modelxx xxxxSettingxx Datasets xxAveragexx
xxAIME 2024  $x$ xxAIME 2025  $x$
Avg@32 Pass@32 Avg@32 Pass@32 Avg@32 Pass@32
Qwen2.5-7B-Instruct Base 11.87 40.00 7.08 33.33 9.48 36.67
AlphaApollo (ours) 11.15 46.67 9.06 46.67 10.11 46.67
[…]
Qwen2.5-14B-Instruct Base 14.17 43.33 13.23 36.67 13.70 40.00
AlphaApollo (ours) 23.33 66.67 14.37 60.00 18.85 63.34
[…]
Qwen2.5-32B-Instruct Base 18.33 40.00 12.60 43.33 15.47 41.67
AlphaApollo (ours) 23.75 63.33 19.27 53.33 21.51 58.33
[…]
Qwen2.5-72B-Instruct Base 18.23 50.00 13.75 30.00 15.99 40.00
AlphaApollo (ours) 22.81 56.67 14.37 53.33 18.59 55.00
[…]
Qwen3-235B-A22B Base 83.44 90.00 79.90 96.67 81.67 93.34
AlphaApollo (ours) 86.67 96.67 84.58 96.67 85.63 96.67
[…]
Llama3.3-70B-Instruct Base 26.56 40.00 5.10 23.33 15.83 31.67
AlphaApollo (ours) 27.71 70.00 21.77 46.67 24.74 58.34
[…]
(a) Tool-call correctness (left: AIME24, right: AIME25)
(b) Accuracy (left: AIME24, right: AIME25)
Figure 7: (a) Tool-call correctness. Most models use tools correctly in nearly or more than 80% of cases. (b) Comparison between responses with or without tool calls. Note that to eliminate the differences between questions, we only collect the questions where at least one response includes a tool call and at least one response does not.
Analysis.
We analyze our system from two perspectives: the quality of tool calls and the impact of tools on final performance. For the quality of tool calls, we extract tool responses from model outputs and check for tool-call errors, such as execution errors in the computational module, querying the retrieval module for nonexistent packages, or attempts to call nonexistent tools. As shown in Figure 7(a), most models use tools correctly in more than 80% of cases, indicating our framework effectively guides models to generate correct tool calls. For the impact of tools on performance, we divide the model outputs into two types: those with tool calls and those without. For questions with both types of responses, we compute accuracies separately. As shown in Figure 7(b), responses with tool calls consistently outperform those without. Notably, with the assistance of tools, Qwen3-235B-A22B and Llama3.3-70B-Instruct are able to solve problems that they completely fail on without tool support, indicating that tools expand the models’ capability frontier.
@@
Figure 8: Case studies illustrating models’ diverse cognitive behaviors exhibited by AlphaApollo. We present the complete model responses in Appendix B.2.
@@
With the support of AlphaApollo’s professional tools, models exhibit diverse cognitive behaviors in their reasoning processes. These behaviors originate from the pre-training data of the base model, which equips the model with the ability to conduct deliberate reasoning (Gandhi et al., 2025). We summarize these cognitive behaviors in tool-augmented reasoning as follows:
  • Decomposition. The model demonstrates the ability to break down a complex problem into smaller, more manageable sub-problems. This strategy not only reduces cognitive load but also increases the likelihood of solving each component correctly, which in turn contributes to the accuracy of the final solution.
@@
4 Related Work
In this section, we systematically review prior work related to our three key features: tool-integrated reasoning in Section 4.1, multi-model reasoning in Section 4.2, and test-time iteration in Section 4.3. Analogous to the Apollo Program, where diverse experts built specialized tools to iteratively launch the Apollo missions, these three features are essential to our AlphaApollo system.
4.1 Tool-integrated Reasoning
@@
Early tool-integrated methods leverage external tools to enhance FMs reasoning, addressing limitations in knowledge and computation. (1) For knowledge, Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Nakano et al., 2021) integrates knowledge bases to furnish FMs with essential facts for accurate inference, though its efficacy hinges on precise retrieval of contextually relevant information without noise (Gao et al., 2023b). To ensure retrieval reliability, Self-Ask (Press et al., 2022) decomposes queries into sub-questions for targeted retrieval, Self-RAG (Asai et al., 2024) verifies and filters irrelevant chunks, and GraphRAG (Edge et al., 2024) structures knowledge as context graphs to capture relational relevance. (2) For computation, Python has emerged as an effective tool for precise computation. Program-aided Language Models (PAL) (Gao et al., 2023a) and Program of Thoughts (PoT) (Chen et al., 2023b) incorporate Python code generation and execution within the reasoning process for arithmetic and logical tasks, while ViperGPT (Surís et al., 2023) extends this to vision, using code to process images and enrich multimodal reasoning.
@@
Beyond tool-integrated methods, agentic frameworks build flexible, tool-integrated environments that allow FMs to invoke tools dynamically, rather than adhering to predefined calling stages. For instance, SciMaster (Chai et al., 2025a) introduces agentic workflows that support dynamic reasoning via a Python-based tool for computation and retrieval, augmented by test-time scaling through reflection, solution refinement, and answer selection mechanisms to enhance FMs’ ability on complex problems. Similarly, OctoTools (Lu et al., 2025) offers a reliable framework that integrates diverse tools through detailed tool cards describing their functions and utilities; it deploys a query analyzer agent to select a task-specific tool subset based on tool cards. This framework enables efficient tool usage when extensive tools are available, yielding more adaptive tool utilization for complex problems. Additionally, Alita (Qiu et al., 2025) proposes a framework for dynamically generating task-specific tools from code. It also leverages web search to iteratively refine both the reasoning process and the design of tools to optimize the solution for real-world tasks.
@@
While the integration of multiple tools unlocks the reasoning potential of FMs, how to let FMs harness the usage of these tools in solving complex problems remains a significant challenge. To address this, tool-learning frameworks (Qin et al., 2024; Liu et al., 2024; Gao et al., 2024) enhance FMs’ tool utilization through targeted post-training. Notably, several frameworks are tailored for tool-integrated long-horizon reasoning, enabling trainable multi-round interactions with tools. For instance, VerlTool (Jiang et al., 2025), RL-Factory (Chai et al., 2025b), and rStar2-Agent (Shang et al., 2025) employ unified tool managers to create model-friendly tool-use environments, supported by established RL training pipelines via VeRL (Sheng et al., 2024). Although these frameworks effectively integrate RL methods into tool-integrated reasoning, the inherent dynamics of such reasoning—particularly long-horizon planning and the incorporation of external tools—pose significant challenges to stable and efficient optimization. To mitigate this, Verl-Agent (Feng et al., 2025) introduces a step-independent rollout mechanism and customizable memory modules, alleviating inherent long-horizon reasoning difficulties. Whereas SimpleTIR (Xue et al., 2025) detects and filters trajectories featuring “void turns”—instances where reasoning collapses and destabilizes multi-turn agentic training—thereby promoting more robust optimization.
4.2 Multi-model Reasoning
@@
Collaboration coordinates multiple models that work synergistically, with each contributing specialized capabilities toward a shared objective. AutoGen (Wu et al., 2024) provides a framework for multi-agent conversations that power next-generation FM applications, enabling agents to jointly tackle tasks such as coding and question answering. MetaGPT (Hong et al., 2024) employs role-based multi-agent collaboration and emphasizes structured, human-like workflows with standard operating procedures (SOPs), assigning roles such as product manager and engineer to inject domain expertise and improve efficiency in software-development tasks. In addition, HuggingGPT (Shen et al., 2023) leverages FMs to orchestrate Hugging Face models across modalities, integrating vision and speech to handle multimodal tasks.
@@
Debate mechanisms engage multiple models in mutual critique and refinement, often paired with tool use to verify facts and resolve ambiguities. The MAD framework (Du et al., 2024) formalizes multi-round proposal, cross-examination, and revision among independent FMs before a final judgment, improving mathematical and strategic reasoning, reducing hallucinations, and working even with black-box models under task-agnostic prompts. CoA (Li et al., 2024) advances this paradigm with a sparse communication topology that lowers computational cost relative to fully connected settings while preserving reasoning performance in experiments with GPT and Mistral. Complementing these efforts, LLM-Coordination (Agashe et al., 2023) examines multi-agent behavior in pure coordination games, finding that FMs excel at environment comprehension yet underperform on theory-of-mind reasoning.
@@
Multi-agent fine-tuning jointly optimizes role-specialized models so that, as a group, they plan, call tools, and summarize more effectively than a single FM. Early systems fine-tune on agent trajectories (FireAct (Chen et al., 2023a)) or curated agent-instruction data (AgentTuning (Zeng et al., 2023)) to endow general agent abilities, providing a foundation for role-based optimization. An emerging direction explicitly fine-tunes a society of models from a common base using inter-agent data to diversify skills and improve coordination (Subramaniam et al., 2025). Shen et al. (2024) decomposes tool-learning into planner, caller, and summarizer roles, each fine-tuned on sub-tasks, achieving superior performance on ToolBench and surpassing single-FM approaches. AgentFly (Zhou et al., 2025) advances this paradigm by introducing memory-based online reinforcement learning for agent fine-tuning without updating FM weights. These methods demonstrate how multi-agent fine-tuning enhances coordination and tool usage in multi-turn tasks.
4.3 Test-time Iteration

### v2

@@ 3 Experiments
Experiment setup.
We evaluate AlphaApollo across various models, from Qwen2.5-1.5B-Insturct to Qwen2.5-14B-Instruct. We evaluate on AIME24, AIME25, CMIMC, HMMT (Feb and Nov), BRUMO, and SMT. We set a maximum of 4 interaction rounds. Details are in Appendix B, C, and D.
Table 1: Agentic reasoning results (Avg@32/Pass@32 in %). Base is evaluated without training and tools; AlphaApollo is evaluated with tools enabled without training. Bold marks the better results.
Dataset Qwen2.5-3B-Instruct Qwen2.5-7B-Instruct Qwen2.5-14B-Instruct
Base AlphaApollo Base AlphaApollo Base AlphaApollo
AIME24 5.21 / 26.67 5.52 / 30.00 12.19 / 36.67 8.85 / 56.67 13.44 / 46.67 16.98/ 60.00
AIME25 3.23 / 36.67 2.19 / 23.33 8.23 / 36.67 6.15 / 36.67 12.29 / 43.33 11.77 / 46.67
CMIMC25 1.17 / 17.50 3.36 / 30.00 4.02 / 30.00 7.42 / 40.00 4.61 / 27.50 11.48 / 40.00
HMMT25 Feb 0.52 / 10.00 2.60 / 16.67 2.08 / 23.33 6.67 / 33.33 3.23 / 23.33 9.58 / 40.00
HMMT25 Nov 3.02 / 20.00 2.50 / 23.33 5.00 / 23.33 6.04 / 26.67 5.73 / 20.00 7.29 / 23.33
BRUMO25 11.25 / 40.00 8.44 / 46.67 18.23 / 50.00 17.18 / 50.00 22.29 / 43.33 22.60 / 60.00
SMT 2025 8.25 / 32.08 8.43 / 41.51 11.62 / 39.62 9.08 / 41.51 14.15 / 41.51 14.74 / 49.06
Average 4.66 / 26.13 4.72 / 30.22 8.77 / 34.23 8.77 / 40.69 10.82 / 35.10 13.49 / 45.58
[…]
Table 2: Agentic learning results (Avg@32 in %) for AlphaApollo. No-training evaluates AlphaApollo using tools without training. LE (MATH-LightEval (hendrycksmath2021)), LIMR (li2025limr), and DS (DeepScaleR (deepscaler2025)) denote training the model on the corresponding datasets.
Dataset Qwen2.5-1.5B-Instruct Qwen2.5-3B-Instruct Qwen2.5-7B-Instruct
No- training AlphaApollo (training) No- training AlphaApollo (training) No- training AlphaApollo (training)
+LE +LIMR +DS +LE +LIMR +DS +LE +LIMR +DS
AIME24 0.63 3.77 3.74 8.96 5.52 9.14 14.35 20.92 8.85 22.91 19.40 25.50
AIME25 0.73 2.68 15.36 14.67 2.19 13.06 14.14 9.33 6.15 16.58 15.28 17.58
CMIMC25 0.63 5.25 7.78 7.11 3.36 6.32 7.80 10.58 7.42 13.16 14.14 16.97
HMMT25 Feb 1.35 4.47 7.27 6.51 2.60 12.27 12.61 14.07 6.67 13.33 9.21 18.30
HMMT25 Nov 0.83 1.73 5.75 4.93 2.50 6.48 5.29 4.21 6.04 11.21 12.28 13.11
BRUMO25 1.98 8.77 15.96 14.98 8.44 23.22 21.30 21.62 17.18 30.26 27.70 33.10
SMT 2025 1.36 5.83 12.07 10.31 8.43 13.11 14.60 10.72 9.08 16.65 18.00 17.88
Average 1.07 4.64 9.70 9.64 4.72 11.94 12.87 13.35 8.77 17.73 16.57 20.35
[…]
Table 3: Agentic evolution results (accuracy in %). w/o Evolution uses tools without evolution, while w/ Evolution enables agentic evolution with tools.
Datasets Qwen2.5-3B-Instruct Qwen2.5-7B-Instruct Qwen2.5-14B-Instruct
w/o Evolution w/ Evolution w/o Evolution w/ Evolution w/o Evolution w/ Evolution
AIME24 6.67 10.00 11.67 12.50 19.17 23.33
AIME25 7.50 7.50 8.33 11.67 18.33 24.17
CMIMC 25 3.12 6.25 10.00 12.50 15.00 19.38
HMMT25 Feb 3.33 5.00 10.00 10.83 15.83 20.00
HMMT25 Nov 1.67 4.17 7.50 7.50 7.50 12.50
BRUMO25 7.50 12.50 11.67 17.50 26.67 31.67
SMT 2025 7.08 8.49 8.02 8.96 13.21 16.51
Average 5.27 7.70 9.60 11.64 16.53 21.08
[…]
@@
We evaluate AlphaApollo from three aspects: agentic reasoning, agentic learning, and agentic evolution. The corresponding results are summarized in Tables 1, 2, and 3, respectively.
Agentic reasoning with tools enables consistent performance gain across model scales.
In Table 1, Figures 5(b), and 5(c), AlphaApollo consistently outperforms no-tool versions across all models, demonstrating improvements from 4.66% to 4.72% on Qwen2.5-3B-Instruct and from 10.82% to 13.49% on Qwen2.5-14B-Instruct. The observed gains in reasoning capability primarily stem from improved tool-call reliability. In Figure 5(a), AlphaApollo’s model-friendly tool-calling module achieves a success rate exceeding 85% across all evaluated datasets. These demonstrate that AlphaApollo consistently excels in boosting FMs’ capabilities, both in reasoning and tool utilization.
Agentic learning in AlphaApollo substantially improves reasoning across model scales.
As shown in Table 2, training on DeepScaleR consistently yields large gains over tool-augmented baselines without training, improving average accuracy by +8.57 on Qwen2.5-1.5B-Instruct, +8.33 on Qwen2.5-3B-Instruct (MATH-LightEval, from 4.72% to 13.35%), and +11.58 on Qwen2.5-7B-Instruct, demonstrating strong scalability of agentic reasoning improvements. Figures 6 and 7 further show steady performance growth on both training and test sets. We additionally compare full-parameter and LoRA training (Figure 8). Full-parameter optimization exhibits faster learning dynamics, with quicker accuracy gains, reduced entropy loss, and consistently higher final performance, indicating the advantage of full-capacity adaptation for agentic reasoning.
Agentic evolution further amplifies reasoning capability through iterative self-improvement.
In Table 3, enabling evolution (+Evo) improves over the non-evolving baseline (w/o Evo). The strongest gains occur on Qwen2.5-14B-Instruct, where average accuracy increases from 16.53% to 21.08% (+4.55), including large boosts on AIME24 and AIME25. Improvements also persist on other backbones, with average gains of +2.43 on Qwen2.5-3B-Instruct. Figure 9 reveals a stable upward trend across evolving rounds with clear scale-dependent gaps, indicating that agentic evolution reliably enhances reasoning performance and scales the benefit with model capability.
Overall, these results show that AlphaApollo delivers consistent, scalable gains by unifying agentic reasoning, learning, and evolution, enabling reliable tool use and iterative refinement of reasoning.
(a) Tool-call success rate (%) in reasoning with AlphaApollo.
(b) Accuracy (%) w/ and w/o tool calls on Qwen2.5-3B-Instruct.
(c) Accuracy (%) w/ and w/o tool calls on Qwen2.5-14B-Instruct.
Figure 5: Tool-call performance and accuracy enhancements in AlphaApollo using Qwen models.
(a) MATH-LightEval
(b) LIMR
(c) DeepScaleR
Figure 6: Training accuracy curves of experiments in Table 2.
(a) MATH-LightEval
(b) LIMR
(c) DeepScaleR
Figure 7: Validation accuracy curves of experiments in Table 2.
Figure 8: Comparison of training and validation in performance and entropy: Full parameters vs. LoRA on Qwen2.5-14B-Instruct with MATH-LightEval.
Figure 9: Accuracy across successive self-evolution rounds for different model scales. Performance consistently improves with more evolution rounds, demonstrating the effectiveness and stability of the self-evolving mechanism.
@@
Figure 10: Case studies of models’ cognitive behaviors exhibited by AlphaApollo.
@@
Under AlphaApollo, models exhibit diverse cognitive reasoning behaviors, rooted in their pre-training data and enabling deliberate reasoning. We showcase in Figure 10 and present complete case studies in Appendix F.2 while summarizing following agentic reasoning behaviors:
  • Decomposition. The model breaks down a complex problem into smaller, more manageable sub-problems. This strategy not only reduces cognitive load but also increases the likelihood of solving each component correctly, which in turn contributes to the accuracy of the final solution.
@@
4 Conclusion
We presented AlphaApollo, a self-evolving agentic reasoning system that addresses limited long-horizon reasoning capacity and unreliable test-time refinement by orchestrating models and tools. AlphaApollo combines multi-turn agentic reasoning, turn-level agentic learning, and multi-round agentic evolution to enable structured tool use, stable RL optimization, and iterative self-improvement with tool-assisted verifications and long-horizon memory. Across seven math reasoning benchmarks and multiple model scales, AlphaApollo delivers consistent gains through reliable tool use, multi-turn reinforcement learning, and further improvements from test-time evolution.
Acknowledgement
We want to thank Guanyu Jiang for his help in testing the early version of AlphaApollo and Yicheng Zhou for his help in testing the evolution part of AlphaApollo.
Appendix A Related Work
In this section, we systematically review prior work related to our three key features: tool-integrated reasoning in Section A.1, multi-model reasoning in Section A.2, and test-time iteration in Section A.3. Analogous to the Apollo Program, where diverse experts built specialized tools to iteratively launch the Apollo missions, these three features are essential to our AlphaApollo system.
A.1 Tool-integrated Reasoning
@@
Early tool-integrated methods leverage external tools to enhance FMs reasoning, addressing limitations in knowledge and computation. (1) For knowledge, Retrieval-Augmented Generation (RAG) (lewis2020retrieval; nakano2021webgpt) integrates knowledge bases to furnish FMs with essential facts for accurate inference, though its efficacy hinges on precise retrieval of contextually relevant information without noise (gao2023retrieval). To ensure retrieval reliability, Self-Ask (press2022selfask) decomposes queries into sub-questions for targeted retrieval, Self-RAG (asai2024selfrag) verifies and filters irrelevant chunks, and GraphRAG (edge2024local) structures knowledge as context graphs to capture relational relevance. (2) For computation, Python has emerged as an effective tool for precise computation. Program-aided Language Models (PAL) (gao2023pal) and Program of Thoughts (PoT) (chen2023program) incorporate Python code generation and execution within the reasoning process for arithmetic and logical tasks, while ViperGPT (suris2023vipergpt) extends this to vision, using code to process images and enrich multimodal reasoning.
@@
Beyond tool-integrated methods, agentic frameworks build flexible, tool-integrated environments that allow FMs to invoke tools dynamically, rather than adhering to predefined calling stages. For instance, SciMaster (chai2025scimaster) introduces agentic workflows that support dynamic reasoning via a Python-based tool for computation and retrieval, augmented by test-time scaling through reflection, solution refinement, and answer selection mechanisms to enhance FMs’ ability on complex problems. Similarly, OctoTools (pan2025octotools) offers a reliable framework that integrates diverse tools through detailed tool cards describing their functions and utilities; it deploys a query analyzer agent to select a task-specific tool subset based on tool cards. This framework enables efficient tool usage when extensive tools are available, yielding more adaptive tool utilization for complex problems. Additionally, Alita (qiu2025alita) proposes a framework for dynamically generating task-specific tools from code. It also leverages web search to iteratively refine both the reasoning process and the design of tools to optimize the solution for real-world tasks.
@@
While the integration of multiple tools unlocks the reasoning potential of FMs, how to let FMs harness the usage of these tools in solving complex problems remains a significant challenge. To address this, tool-learning frameworks (qin2024toolllm; liu2024apigen; gao2024confucius) enhance FMs’ tool utilization through targeted post-training. Notably, several frameworks are tailored for tool-integrated long-horizon reasoning, enabling trainable multi-round interactions with tools. For instance, VerlTool (jiang2025verltool), RL-Factory (chai2025rlfactory), and rStar2-Agent (shang2025rstar2) employ unified tool managers to create model-friendly tool-use environments, supported by established RL training pipelines via VeRL (sheng2024hybridflow). Although these frameworks effectively integrate RL methods into tool-integrated reasoning, the inherent dynamics of such reasoning—particularly long-horizon planning and the incorporation of external tools—pose significant challenges to stable and efficient optimization. To mitigate this, Verl-Agent (feng2025group) introduces a step-independent rollout mechanism and customizable memory modules, alleviating inherent long-horizon reasoning difficulties. Whereas SimpleTIR (xue2025simpletir) detects and filters trajectories featuring “void turns”—instances where reasoning collapses and destabilizes multi-turn agentic training—thereby promoting more robust optimization.
A.2 Multi-model Reasoning
@@
Collaboration coordinates multiple models that work synergistically, with each contributing specialized capabilities toward a shared objective. AutoGen (wu2024autogen) provides a framework for multi-agent conversations that power next-generation FM applications, enabling agents to jointly tackle tasks such as coding and question answering. MetaGPT (hong2024metagpt) employs role-based multi-agent collaboration and emphasizes structured, human-like workflows with standard operating procedures (SOPs), assigning roles such as product manager and engineer to inject domain expertise and improve efficiency in software-development tasks. In addition, HuggingGPT (shen2023hugginggpt) leverages FMs to orchestrate Hugging Face models across modalities, integrating vision and speech to handle multimodal tasks.
@@
Debate mechanisms engage multiple models in mutual critique and refinement, often paired with tool use to verify facts and resolve ambiguities. The MAD framework (du2024improving) formalizes multi-round proposal, cross-examination, and revision among independent FMs before a final judgment, improving mathematical and strategic reasoning, reducing hallucinations, and working even with black-box models under task-agnostic prompts. CoA (li2024improving) advances this paradigm with a sparse communication topology that lowers computational cost relative to fully connected settings while preserving reasoning performance in experiments with GPT and Mistral. Complementing these efforts, LLM-Coordination (agashe2023llm) examines multi-agent behavior in pure coordination games, finding that FMs excel at environment comprehension yet underperform on theory-of-mind reasoning.
@@
Multi-agent fine-tuning jointly optimizes role-specialized models so that, as a group, they plan, call tools, and summarize more effectively than a single FM. Early systems fine-tune on agent trajectories (FireAct (chen2023fireact)) or curated agent-instruction data (AgentTuning (zeng2023agenttuning)) to endow general agent abilities, providing a foundation for role-based optimization. An emerging direction explicitly fine-tunes a society of models from a common base using inter-agent data to diversify skills and improve coordination (subramaniam2025multiagentft). shen2024small decomposes tool-learning into planner, caller, and summarizer roles, each fine-tuned on sub-tasks, achieving superior performance on ToolBench and surpassing single-FM approaches. AgentFly (zhou2025agentfly) advances this paradigm by introducing memory-based online reinforcement learning for agent fine-tuning without updating FM weights. These methods demonstrate how multi-agent fine-tuning enhances coordination and tool usage in multi-turn tasks.
A.3 Test-time Iteration